A triangle has three angles A, B,C. Angle A is 38 less than three times angle C, and angle B is 14 more than twice angle C. Find the measures of all the angles.
Answer by mananth(16946)

let angle c be x deg
Angle b =2x+14
Angle a= 3x-38
...
x+2x+14+3x-38=180
6x-24=180
6x=204
/6
x=34 deg. Angle C
Angle b = 2x+14 = 68+14=82 deg
Angle a =3x-38=102-38=64 deg

...
m.ananth@hotmail.ca